# SoK: A Literature and Engineering Review of Regular Expression Denial of Service (ReDoS)

### v3

@@ Abstract.
Regular Expression Denial of Service (ReDoS) is a vulnerability class that has become prominent in recent years. Attackers can weaponize such weaknesses as part of asymmetric cyberattacks that exploit the slow worst-case matching time of regular expression (regex) engines. In the past, problematic regular expressions have led to outages at Cloudflare and Stack Overflow, showing the severity of the problem. While ReDoS has drawn significant research attention, there has been no systematization of knowledge to delineate the state of the art and identify opportunities for further research. In this paper, we describe the existing knowledge on ReDoS. We first provide a systematic literature review, discussing approaches for detecting, preventing, and mitigating ReDoS vulnerabilities. Then, our engineering review surveys the latest regex engines to examine whether and how ReDoS defenses have been realized. Combining our findings, we observe that (1) in the literature, almost no studies evaluate whether and how ReDoS vulnerabilities can be weaponized against real systems, making it difficult to assess their real-world impact; and (2) from an engineering view, many mainstream regex engines now have ReDoS defenses, rendering many threat models obsolete. We conclude with an extensive discussion, highlighting avenues for future work. The open challenges in ReDoS research are to evaluate emerging defenses and support engineers in migrating to defended engines. We also highlight the parallel between performance bugs and asymmetric DoS, and we argue that future work should capitalize more on this similarity and adopt a more systematic view on ReDoS-like vulnerabilities.
Systematization of Knowledge (SoK), Regular Expression Denial of Service (ReDoS), Regex Engines, ReDoS Mitigation Strategies
Both authors contributed equally to this research.
ccs: General and reference Surveys and overviewsccs: Security and privacy Denial-of-service attacksccs: Software and its engineering Software libraries and repositoriesconference: ; ;
@@ 1. Introduction
Regular Expression Denial of Service (ReDoS) (Crosby, 2003; Goyvaerts, 2003; Roichman and Weidman, 2009) is a type of denial-of-service (DoS) attack (Gligor, 1983; noa, [n. d.]) that exploits inefficiencies in regular expression (regex) engines. Many regex engines have worst-case exponential backtracking behavior. By crafting input strings that trigger this behavior, attackers can cause systems to consume disproportionate CPU resources, leading to service disruptions. ReDoS threatens real systems: slow regex matches caused outages at Stack Overflow (stackstatus, 2016) and Cloudflare (clo, 2019). ReDoS is the fourth most common vulnerability in the JavaScript npm package ecosystem (Bhuiyan et al., 2023) and one of the fastest-growing vulnerability classes within the same ecosystem (sny, 2019), triggering more than one million pull requests (Section D.5). Consequently, there has been much research on ReDoS, with dozens of papers since 2015 (Section 3). However, as yet there has been no systematization of knowledge about ReDoS, making it challenging to consolidate existing knowledge, identify gaps, and guide future efforts to address this complex and evolving threat.
@@
To address this gap, we conducted a comprehensive literature review of recent papers on ReDoS (Section 3). Our goals were to provide a tutorial on regular expressions and ReDoS, summarize previous work, assess progress in the field, and identify future work opportunities. We analyze 35 papers on ReDoS vulnerabilities from top security conferences, focusing on detection, prevention, and mitigation strategies, to provide a comprehensive overview of the current state of the field. One primary observation from our review was that almost all prior works rely on at least one of two main models of regex engine behavior: the first introduced in 1968 by Thompson (Thompson, 1968), and the second in 1994 by Spencer (Spencer, 1994). While these models have been instrumental in shaping foundational research, they do not reflect modern regex engines, many of which now incorporate defenses against ReDoS attacks. ReDoS research must be situated within real regex engine implementations, motivating us to perform an updated engineering review.
@@
Thus, we conducted an engineering review of modern regex engines to address this disconnect between academic research and practical applications and update the understanding of the capabilities of regex engines (Section 4). This review examines the implementations in the major programming languages to evaluate how they handle ReDoS vulnerabilities, either through new algorithms, resource caps, or other mitigations. By analyzing these engines’ designs, defenses, and real-world adoption, we provide a detailed account of the current state of regex processing and its implications for ReDoS research. This engineering perspective complements the findings of our literature review and highlights the differences between traditional regex engines and contemporary regex engine behavior, offering up-to-date insights for future studies.
@@
As a result of our synthesis of academic knowledge and engineering analysis, we report five findings from this systematization of knowledge. In our review of the academic literature, we identified areas where threat models, ReDoS definitions, and attack evaluations could benefit from further refinement and alignment. In our engineering review of the regex engines, we note that four major engines have used three distinct algorithms to mitigate ReDoS, but these defenses do not always reduce the time complexity to linear. We encourage the research community to explore opportunities to improve conceptual clarity in future ReDoS research and to focus efforts on evaluating and advancing state-of-the-art ReDoS defenses. To support researchers, we propose a framework described in Appendix C, which provides a uniform evaluation environment and also allows comparisons between different regex engines.
@@ 2.1.2. Extended regexes (E-regexes)
Over time, to simplify the representation, the K-regexes were “extended” (E-regexes) with additional operators and constructs. Syntax sugar notations, like character ranges (e.g., [A-Z] for A|B|...|Z) and one-or-more-repetition (R+), are still regular and can be transformed into equivalent K-regexes. K-regexes were also enhanced (Friedl, 2006; Hazel, 1997) with irregular features such as backreferences and zero-width assertions (Aho, 1980).
Yet, irregular features increase expressive power but exceed the boundaries of regular languages (i.e., they cannot be directly modeled with finite automata), complicating theoretical analysis, implementation, and matching performance. A complete set of regex notation is given in Appendix A.
@@ 2.2. Regex Engines
Regexes are supported in all mainstream programming languages (Friedl, 2006) and are part of many other systems, e.g., intrusion detection and prevention systems (Roesch, 1999; Reg, [n. d.]), web application firewalls (Waf, [n. d.]), and logging tools (Buchanan, 2023; Acc, [n. d.]). These systems all implement or integrate a component called a regex engine that provides support for the regex tasks of recognition and parsing. As summarized by Davis (Davis, 2020b), typical regex engine implementations followed one of two algorithms. The first is a depth-first, backtracking search based on a library by Spencer (Spencer, 1994). The second is a breadth-first, “lockstep” search proposed by Thompson (Thompson, 1968)
The characteristics of these engine styles are summarized in Table 1. For the simple case of K-regexes, both can be modeled as NFA simulations. [^1] Implementations of the Thompson-style typically have a direct mapping to an NFA simulation, while implementations of the Spencer-style may be implemented as recursive descent parsers. Here, we treat K-regexes, where both amount to NFA simulations. These approaches differ in their handling of ambiguity (Allauzen et al., 2008)—where multiple transitions in the NFA are possible.
@@
Table 1. Tradeoffs of Spencer and Thompson algorithms for regex engines (Cox, 2007; Davis, 2020b). Thompson’s algorithm uses an efficient algorithm over an automata representation to obtain linear time complexity (in input string length), but this formal approach reduces its expressiveness. Spencer gains E-regex support at the cost of exponential match times.
@@
Most of the time (for most regexes, for most inputs, in most application contexts) both approaches run quickly enough, and do not form a bottleneck for the broader application. However, while the Thompson approach guarantees match times that are linear in the input length, the Spencer approach exhibits worst-case polynomial or exponential match times on certain regex-input combinations. Figure 2 gives two examples where Spencer’s algorithm exhibits polynomial and exponential time complexity, respectively. In each case, on an input $w=aa...ab$, the regex engine encounters ambiguity when processing an “a” symbol. At the ultimate mismatch (caused by the final “b”), the backtracking will take $|w|^{2}$ time for example (a) and $2^{|w|}$ time for example (b). This worst-case behavior is colloquially termed “catastrophic backtracking.” Note that the actual cost of the match depends on the length of the input; the term pumping refers to the repetition of the problematic substring of the input (each “a” in this example).
$q_{0}$$q_{1}$$\epsilon$$a$$a$
@@
Figure 2. Example regexes that, in Spencer’s algorithm, exhibit time complexity that is either: (a) polynomial or (b) exponential; in the length of the input string “aa…ab”.
@@
2.3. Regex Denial of Service (ReDoS)
@@ 2.3.1. ReDoS in Principle
This engineering tradeoff—prioritizing expressiveness and accepting poor worst-case performance—has led to a security vulnerability called ReDoS (Crosby, 2003). ReDoS is assigned CWE-1333: Inefficient Regular Expression Complexity (MITRE, 2024). ReDoS is an example of an algorithmic complexity attack (Crosby and Wallach, 2003), characterized by the small cost to create a malicious input and the large cost to process it. Algorithmic complexity attacks are one kind of asymmetric DoS attack (Mantas et al., 2015), where the asymmetry comes from the worst-case performance of the data structures and algorithms used to implement the system.
@@
    A Backtracking Regex Engine: The regex engine employed for pattern matching in the victim system utilizes a backtracking approach, as discussed in Section 2.2. Many regex engines implemented in programming languages (e.g., JavaScript, Java, Python) have at some point met this criterion (Section 4).
@@
    A Vulnerable Regex: The victim system uses a super-linear (SL), problematically ambiguous regex; for which the attacker crafts attack strings that take polynomial or exponential time to match in the string length (Wüstholz et al., 2017; Weideman et al., 2016).
@@
Given these three ingredients, if the system has insufficient mitigations (safeguards) in place, then the attacker can cause the victim system to consume excessive computational resources (CPU time). The attacker may need to tailor the number of pumps based on the length of typical input (to avoid anomaly detection) or the maximum length allowed (to avoid truncation).
Figure 3. Distribution of OWASP Top-10 and ReDoS CVEs per ecosystem, showing average prevalence per category. Data cover 2014-2023. For full methodology, see Appendix B.
@@ 2.3.2. ReDoS in Practice
In recent years, ReDoS has become a common security concern. Two major web services had outages that resulted from slow regex behavior: Stack Overflow in 2016 (stackstatus, 2016) and Cloudflare in 2019 (clo, 2019). While these outages did not involve an adversary (i.e., they were reliability issues, not security issues), they made service providers aware of the risk of ReDoS. Since 2020, the security company Snyk has observed a jump in reported ReDoS vulnerabilities, especially in the JavaScript package ecosystem npm (sny, 2019). Grey literature resources such as GitHub’s security team give advice for repairing ReDoS vulnerabilities (sny, [n. d.]; Backhouse, 2023; Davis, 2020a; reg, [n. d.]).
For a sense of the formal disclosure of ReDoS vulnerabilities, we examined the National Vulnerability Database for CVE records pertaining to ReDoS. Figure 3 shows the result. Following the keyword method of Hassan et al. (Hassan et al., 2023), since 2014 $\sim$400 ReDoS CVEs have been disclosed in software from six major programming language ecosystems. As a reference point, we compared these results to CVEs related to the “OWASP Top-10”. Averaging over 2014-2023, the rate of ReDoS exceeds four of the OWASP Top-10 vulnerability types, though it is far below the Top-5.
@@ 3. Academic Research on ReDoS
  • Exclusion Criteria: Papers that focus on general denial of service or those that discuss regular expressions, without specifically addressing ReDoS.
@@
    Vulnerability Definitions: Focuses on how papers describe what constitutes a ReDoS vulnerability or a vulnerable regex, e.g., having exponential matching time.
@@
We begin by discussing the main existing research directions (Section 3.1), distinguishing between detection, prevention, mitigation, and other related works. We then explore the assumptions made about the attackers’ capabilities or the underlying systems (Section 3.2), along with the definitions of ReDoS vulnerabilities or successful attacks provided in prior works (Section 3.3). Finally, we discuss how the studied papers were evaluated (Section 3.4).
@@ 3.1. Main Existing Research Directions
Figure 4 provides an overview of the main research directions in the study of ReDoS vulnerabilities, categorized into detection, prevention, mitigation, and complementary studies. Detection focuses on identifying ReDoS vulnerabilities in regular expressions and understanding their exploitability. Prevention aims to address the root causes of ReDoS vulnerabilities by improving regex engines or restricting vulnerable constructs. Mitigation, on the other hand, focuses on minimizing the impact of ReDoS attacks by deploying runtime safeguards. We discuss these categories in detail below.
@@ 3.1.1. Detection
Detection is a critical area of research in ReDoS, with studies focusing on identifying vulnerable regexes across various ecosystems. Nine papers analyze open-source libraries, such as npm (Bhuiyan et al., 2023; Staicu and Pradel, 2018), pip (Davis et al., 2018a; Wang et al., 2023), and Maven (Wang et al., 2023; Li et al., 2021), while seven focus on testing live web applications (Staicu and Pradel, 2018; Barlas et al., 2022). Additionally, 27 papers investigate ReDoS vulnerabilities using curated regex datasets (Turoňová et al., 2022; Liu et al., 2021; Siddiq et al., 2024a, b). We categorize detection methods into five key approaches: Heuristic-Based Detection (three papers), Automata-Based Detection (seven papers), Dynamic Detection (four papers), Hybrid Detection (five papers), and Machine Learning-Based Detection (one paper). Heuristic-based techniques rely on simple pattern rules to flag problematic regexes, offering speed but often at the cost of high false positive rates. Automata-based approaches, which model regexes using finite-state machines, provide formal accuracy but are computationally expensive. Dynamic detection evaluates regex behavior on specific inputs, revealing real-world vulnerabilities but requiring significant runtime resources. Hybrid methods combine static and dynamic techniques to balance precision and efficiency, while emerging machine learning approaches use regex datasets to predict vulnerabilities, opening new avenues but facing challenges like adversarial examples and limited training data. This categorization highlights the trade-offs in speed, accuracy, and false positive/negative rates among different techniques, offering a structured view of existing detection strategies and their potential for future enhancement.
@@
Heuristic-Based Detection (3 studies) There are many approaches that employ simple pattern-matching techniques to identify potentially vulnerable regular expressions. Studies by Kluban et al. (Kluban et al., 2024, 2022) and Davis et al. (Davis et al., 2017) utilize tools like regexploit (noa, 2024b), redos-detector (jen, 2024), and safe-regex (noa, 2024a) to identify these vulnerabilities by searching for constructs such as infinite repeats (a*b*a*), branches ($a|b$), or nested quantifiers (*, +, ?, {n,m}), which are often associated with ReDoS vulnerabilities. While these methods are fast, enabling the rapid analysis of large numbers of regular expressions, they typically focus on syntax rather than semantics, leading to potential accuracy issues. For instance, Parolini et al. (Parolini and Miné, 2023) demonstrated that while tools like regexploit achieved low false positive and negative rates, others like safe-regex and redos-detector exhibited significantly higher false positive rates. This highlights the primary limitation of heuristic-based approaches, which often require further analysis to confirm vulnerabilities.
@@
Automata-based Static Analysis (6 studies) Automata-based static analysis is a powerful approach for detecting problematic regular expressions without requiring program execution. This method has been explored in several studies (Rathnayake and Thielecke, 2017; Kirrage et al., 2013; Parolini and Miné, 2023; Turoňová et al., 2022; Su et al., 2024; Weideman et al., 2016), which employ automata models such as NFAs, DFAs, and enhanced automata to evaluate regex semantics and detect vulnerabilities that lead to ReDoS attacks.

### v4

@@ Abstract.
Regular Expression Denial of Service (ReDoS) is a vulnerability class that has become prominent in recent years. Attackers can weaponize such weaknesses as part of asymmetric cyberattacks that exploit the slow worst-case matching time of regular expression (regex) engines. In the past, problematic regexes have led to outages at Cloudflare and Stack Overflow, showing the severity of the problem. While ReDoS has drawn significant research attention, there has been no systematization of knowledge to delineate the state of the art and identify opportunities for further research. In this paper, we describe the existing knowledge on ReDoS. We first provide a systematic literature review, discussing approaches for detecting, preventing, and mitigating ReDoS vulnerabilities. Then, our engineering review surveys the latest regex engines to examine whether and how ReDoS defenses have been realized. Combining our findings, we observe that (1) in the literature, almost no studies evaluate whether and how ReDoS vulnerabilities can be weaponized against real systems, making it difficult to assess their real-world impact; and (2) from an engineering view, many mainstream regex engines have introduced partial or full ReDoS defenses, rendering many threat models obsolete. We conclude by highlighting avenues for future work. The open challenges in ReDoS research are to evaluate emerging defenses and support engineers in migrating to defended engines. We also highlight the parallel between performance bugs and asymmetric DoS, and we argue that future work should capitalize more on this similarity and adopt a more systematic view on ReDoS-like vulnerabilities.
Systematization of knowledge (SoK), regular expression denial of service (ReDoS), regex engines, ReDoS defenses
conference: ACM Asia Conference on Computer and Communications Security; August 25–29, 2025; Hanoi, Vietnambooktitle: ACM Asia Conference on Computer and Communications Security (ASIA CCS ’25), August 25–29, 2025, Hanoi, Vietnamcopyright: ccjournalyear: 2025doi: 10.1145/3708821.3733912isbn: 979-8-4007-1410-8/2025/08ccs: General and reference Surveys and overviewsccs: Security and privacy Denial-of-service attacksccs: Software and its engineering Empirical software validation
@@ 1. Introduction
Regular Expression Denial of Service (ReDoS) (Crosby, 2003; Goyvaerts, 2021; Roichman and Weidman, 2009) is a type of denial of service (DoS) attack (Gligor, 1983; Nsrav, 2024) that exploits inefficiencies in regular expression (regex) engines. Many regex engines have worst-case exponential backtracking behavior. By crafting input strings that trigger this behavior, attackers can cause systems to consume disproportionate CPU resources, leading to service disruptions. ReDoS threatens real systems: slow regex matches caused outages at Stack Overflow (Status, 2016) and Cloudflare (Graham-Cumming, 2019). ReDoS is the fourth most common server-side vulnerability class in JavaScript’s npm ecosystem—only path traversal, prototype pollution, and command injection appear more often (Bhuiyan et al., 2023). It is also among the fastest-growing vulnerabilities in the same ecosystem (Limited, 2019). Consequently, there has been much research on ReDoS, with dozens of papers since 2015 (§3). However, as yet there has been no systematization of knowledge about ReDoS, making it challenging to consolidate existing knowledge, identify gaps, and guide future efforts to address this complex and evolving threat.
@@
To address this gap, we conducted a comprehensive literature review of recent papers on ReDoS (§3). Our goals were to provide a tutorial on regexes and ReDoS, summarize previous work, assess progress in the field, and identify future work opportunities. We analyze 35 papers on ReDoS vulnerabilities from top security conferences, focusing on detection, prevention, and mitigation strategies, to provide a comprehensive overview of the current state of the field. One primary observation from our review was that almost all prior works rely on at least one of two main models of regex engine behavior: the first introduced in 1968 by Thompson (Thompson, 1968), and the second in 1994 by Spencer (Spencer, 1994). While these models have been instrumental in shaping foundational research, they do not fully reflect modern regex engines, many of which now incorporate defenses against ReDoS attacks. ReDoS research must be situated within real regex engine implementations, motivating us to perform an updated engineering review.
@@
Thus, we conducted an engineering review of modern regex engines to address this disconnect between academic research and practical applications and update the understanding of the capabilities of regex engines (§4). This review examines the implementations in the major programming languages to evaluate how they handle ReDoS vulnerabilities, either through new algorithms, resource caps, or other mitigations. By analyzing these engines’ designs, defenses, and real-world adoption, we provide a detailed account of the current state of regex processing and its implications for ReDoS research. This engineering perspective complements the findings of our literature review and highlights the differences between traditional regex engines and contemporary regex engine behavior, offering up-to-date insights for future studies.
@@
As a result of our synthesis of academic knowledge and engineering analysis, we discuss five main findings (§5). In our review of the academic literature, we identified areas where threat models, ReDoS definitions, and attack evaluations could benefit from further refinement and alignment. In our engineering review of regex engines, we note that four major engines have used three distinct algorithms to mitigate ReDoS, but these defenses do not always reduce the time complexity to linear. We encourage the research community to explore opportunities to improve conceptual clarity in future ReDoS research and to focus efforts on evaluating and advancing state-of-the-art ReDoS defenses.
@@ 2.1.2. Extended regexes (E-regexes)
Over time, to simplify the representation, the K-regexes were “extended” (E-regexes) with additional operators and constructs. Syntax sugar notations, like character ranges (e.g., [A-Z] for A|B|...|Z) and one-or-more-repetition (R+), are still regular and can be transformed into equivalent K-regexes. Some more complex extensions, such as atomic groups (Berglund et al., 2017) and lookarounds (e.g.(Chattopadhyay et al., 2025; Berglund et al., 2021; Fujinami and Hasuo, 2024)), have also been shown to be regular.
Many other extensions have been made (Friedl, 2006; Hazel, 1997), including irregular features such as backreferences (Aho, 1980). These extensions increase expressive power but exceed the boundaries of regular languages (i.e., they cannot be directly modeled with finite automata), complicating theoretical analysis, implementation, and matching performance. A curated set of regex features is given in Appendix A.
@@ 2.2. Regex Engines
Regexes are supported in all mainstream programming languages (Friedl, 2006) and are part of many other systems (e.g., intrusion detection and prevention systems (Roesch, 1999; Team, 2024a), web application firewalls (Amazon Web Services, 2024), and logging tools (Buchanan, 2023; B.V., 2024)). These systems all implement or integrate a component called a regex engine that provides support for the regex tasks of recognition and parsing. As summarized by Davis (Davis, 2020b), typical regex engine implementations followed one of two algorithms. The first is a depth-first, backtracking search based on a library by Spencer (Spencer, 1994). The second is a breadth-first, lockstep search proposed by Thompson (Thompson, 1968).
The characteristics of these engine styles are summarized in Table 1. For the simple case of K-regexes, both can be modeled as NFA simulations. [^1] Implementations of Thompson’s algorithm typically have a direct mapping to an NFA simulation, while implementations of Spencer’s algorithm may be implemented as recursive descent parsers. Here, we treat K-regexes, where both amount to NFA simulations. These approaches differ in their handling of ambiguity (Allauzen et al., 2008)—where multiple transitions in the NFA are possible.
@@
Table 1. Tradeoffs of Spencer and Thompson algorithms for regex engines (Cox, 2007; Davis, 2020b). Thompson’s algorithm uses an efficient algorithm over an automata representation to obtain linear time complexity (in input string length), but this formal approach reduces its expressiveness. Spencer’s algorithm gains broader E-regex support at the cost of exponential match times.
@@
Most of the time (for most regexes, inputs, and application contexts), both approaches run quickly enough and do not form a bottleneck for the broader application. However, while the Thompson’s algorithm approach guarantees match times that are linear in the input length, Spencer’s algorithm exhibits worst-case polynomial or exponential match times on certain regex-input combinations. Figure 2 gives two examples where Spencer’s algorithm exhibits polynomial and exponential time complexity, respectively. In each case, on an input w=aaabw=$a$...ab, the regex engine encounters ambiguity when processing an “a” symbol. At the ultimate mismatch (caused by the final “b”), the backtracking will take $|w|^{2}$ time for example (a) and $2^{|w|}$ time for example (b). This worst-case behavior is colloquially termed catastrophic backtracking. Note that the actual cost of the match depends on the length of the input; the term pumping refers to the repetition of the problematic substring of the input (each “a” in this example).
[…]
@@
[…]
2.3. Regular Expression Denial of Service (ReDoS)
@@ 2.3.1. ReDoS in Principle
Regex engines often favor expressive power over predictable worst-case efficiency, and that engineering trade-off leaves them open to ReDoS (Crosby, 2003). The Common Weakness Enumeration catalog labels the problem CWE-1333: Inefficient Regular Expression Complexity (MITRE, 2021). ReDoS is an example of an algorithmic complexity attack (Crosby and Wallach, 2003), characterized by the small cost to create a malicious input and the large cost to process it. Algorithmic complexity attacks are one kind of asymmetric DoS attacks (Mantas et al., 2015), where the asymmetry comes from the worst-case performance of the data structures and algorithms used to implement the system.
@@
    A Backtracking Regex Engine: The regex engine employed for pattern matching in the victim system utilizes a backtracking approach (§2.2). Many regex engines implemented in programming languages (e.g., JavaScript, Java, Python) have at some point met this criterion (§4).
@@
    A Vulnerable Regex: The victim system uses a super-linear, problematically ambiguous regex; for which the attacker crafts attack strings that take polynomial or exponential time to match in the string length (Wüstholz et al., 2017; Weideman et al., 2016).
@@
Given these three ingredients, if the system has insufficient mitigations (safeguards) in place, then the attacker can cause the victim system to consume excessive computational resources (e.g., CPU and memory). The attacker may need to tailor the number of pumps based on the length of typical input (to avoid anomaly detection) or the maximum length allowed (to avoid truncation).
Figure 3. Distribution of OWASP Top 10 and ReDoS CVEs per ecosystem, showing average prevalence per category. N indicates the total number of CVEs recorded for each ecosystem. The data cover 2014-2023. For details, see Appendix B.
@@ 2.3.2. ReDoS in Practice
In recent years, ReDoS has become a common security concern. Two major web services had outages that resulted from slow regex behavior: Stack Overflow in 2016 (Status, 2016) and Cloudflare in 2019 (Graham-Cumming, 2019). While these outages did not involve an adversary (i.e., they were reliability issues, not security issues), they made service providers aware of the risk of ReDoS. Since 2020, the security company Snyk has observed a jump in reported ReDoS vulnerabilities, especially in the JavaScript package ecosystem npm (Limited, 2019). In parallel with this surge, a rich body of grey literature has emerged to offer practical guidance on repairing ReDoS vulnerabilities (Limited, 2024; Backhouse, 2023; Davis, 2020a; Goyvaerts, 2019).
For a sense of the formal disclosure of ReDoS vulnerabilities, we examined the National Vulnerability Database (NVD) for Common Vulnerabilities and Exposures (CVE) records pertaining to ReDoS. Following the keyword method of Hassan et al. (Hassan et al., 2023), since 2014 $\sim$400 ReDoS CVEs have been disclosed in software from six major programming language ecosystems. As a reference point, we compared these results to CVEs related to the OWASP Top 10 (Foundation, 2021). Averaging over 2014-2023, Figure 3 shows that the rate of ReDoS exceeds four of the OWASP Top 10 vulnerability types.
@@ 3. Academic Research on ReDoS
  • Exclusion Criteria: Papers that focus on general DoS or those that discuss regexes, without specifically addressing ReDoS. Appendix C provides further discussion of the excluded works.
@@
    Vulnerability Definitions: Focuses on how papers describe what constitutes a ReDoS vulnerability or a vulnerable regex (e.g., having exponential matching time).
@@
We begin by discussing the main existing research directions (§3.1), distinguishing between detection, prevention, mitigation, and other related works. We then explore the assumptions made about the attackers’ capabilities or the underlying systems (§3.2), along with the definitions of ReDoS vulnerabilities or successful attacks provided in prior works (§3.3). Finally, we discuss how the studied papers were evaluated (§3.4).
@@ 3.1. Main Existing Research Directions
Figure 4 provides an overview of the main research directions in the study of ReDoS vulnerabilities, categorized into detection, prevention, mitigation, and complementary studies. Detection focuses on identifying ReDoS vulnerabilities in regexes and understanding their exploitability. Prevention aims to address the root causes of ReDoS vulnerabilities by improving regex engines or restricting vulnerable constructs. Mitigation, on the other hand, focuses on minimizing the impact of ReDoS attacks by deploying runtime safeguards. We discuss these categories in detail below.
@@ 3.1.1. Detection
Detection is a critical area of research in ReDoS, with studies focusing on identifying vulnerable regexes across various ecosystems. Nine papers analyze open-source package ecosystems, such as npm (Bhuiyan et al., 2023; Staicu and Pradel, 2018), pip (Davis et al., 2018a; Wang et al., 2023), and Maven (Wang et al., 2023; Li et al., 2021a), while seven focus on testing live web applications (Staicu and Pradel, 2018; Barlas et al., 2022). Additionally, 27 papers investigate ReDoS vulnerabilities using curated regex datasets (Turoňová et al., 2022; Liu et al., 2021; Siddiq et al., 2024a, b). We categorize detection methods into five key approaches: Heuristic-based Detection (three papers), Automata-based Detection (seven papers), Dynamic Detection (four papers), Hybrid Detection (five papers), and Machine Learning (ML)-based Detection (one paper). Heuristic-based techniques rely on simple pattern rules to flag problematic regexes, offering speed but often at the cost of high false positive rates. Automata-based approaches, which model regexes using finite-state machines, provide formal accuracy but are computationally expensive. Dynamic detection evaluates regex behavior on specific inputs, revealing real-world vulnerabilities but requiring significant runtime resources. Hybrid methods combine static and dynamic techniques to balance precision and efficiency, while emerging ML approaches use regex datasets to predict vulnerabilities, opening new avenues but facing challenges like adversarial examples and limited training data. This categorization highlights the trade-offs in speed, accuracy, and false positive/negative rates among different techniques, offering a structured view of existing detection strategies and their potential for future enhancement.
@@
Heuristic-based Detection (3 studies): There are many approaches that employ simple pattern-matching techniques to identify potentially vulnerable regexes. Studies by Kluban et al. (Kluban et al., 2024, 2022) and Davis et al. (Davis et al., 2017) utilize tools like regexploit (LLC., 2024), redos-detector (Jenkinson, 2024), and safe-regex (Davis, 2024a) to identify these vulnerabilities by searching for constructs such as infinite repeats (a*b*a*), branches ($a|b$), or nested quantifiers (*, +, ?, {n,m}), which are often associated with ReDoS vulnerabilities. While these methods are fast, enabling the rapid analysis of large numbers of regexes, they typically focus on syntax rather than semantics, leading to potential accuracy issues. For instance, Parolini et al. (Parolini and Miné, 2023) demonstrated that while tools like regexploit achieved low false positive and negative rates, others like safe-regex and redos-detector exhibited significantly higher false positive rates. This highlights the primary limitation of heuristic-based approaches, which often require further analysis to confirm vulnerabilities.
Automata-based Static Analysis (6 studies):
@@
Automata-based static analysis is a powerful approach for detecting problematic regexes without requiring program execution. This method has been explored in several studies (Rathnayake and Thielecke, 2017; Kirrage et al., 2013; Parolini and Miné, 2023; Turoňová et al., 2022; Su et al., 2024; Weideman et al., 2016), which employ automata models such as NFAs, DFAs, and enhanced automata to evaluate regex semantics and detect vulnerabilities that lead to ReDoS attacks.